Marme Inc has preferred stock selling for 96 percent of par that pays an annual coupon of 10 percent

What would be Marmes component cost of preferred stock?
Answer #1
Cost of preferred stock = Annual coupon / Current Price
= 10 / 96
= 10.42%
Working:
Suppose par value is 100.
Annual coupon = Par value * Coupon rate
= 100*10%
= 10
Current Price = Par Value * % of Par Value
= 100*96%
= 96